OpenStudy (anonymous):

I have another one m/4.5=2/5

jigglypuff314 (jigglypuff314):

multiply both sides by 4.5

OpenStudy (anonymous):

what do you mean multiply by both sides? I am not very good at algebra sorry

jigglypuff314 (jigglypuff314):

\[\frac{ m }{ 4.5 }= \frac{ 2 }{ 5 } \rightarrow \frac{ m }{ 4.5 }\times 4.5= \frac{ 2 }{ 5 }\times 4.5 \rightarrow m = \frac{ 2 }{ 5 }\times 4.5\]

OpenStudy (anonymous):

ok so the results is ?

OpenStudy (anonymous):

m=1.5?

jigglypuff314 (jigglypuff314):

m = 1.8

OpenStudy (anonymous):

thanks